A sump pump (used to drain water from the basement of houses built below the water table) is draining a flooded basement at the rate of 0.750 L/s, with an output pressure of 3.00×105 N/m2.


  1. The water enters a hose with a 3.00-cm inside diameter and rises 2.50 m above the pump. What is its pressure at this point? Blank 1
  2. The hose goes over the foundation wall, losing 0.500 m in height, and widens to 4.00 cm in diameter. What is the pressure now? Blank 2

Round your answer to the nearest 0 decimal place

1
Expert's answer
2021-05-04T06:40:18-0400

Answer :-

from Bernoulli's Equation


"p+\frac{1}{2}\rho v^2 +\rho gh = cons."



"P \ is \ pressure \ of \ the \ fluid \\
\rho \ is \ the \ densty \ of \ the \ fluid \\
g \ is \ the \ acceleration \ due \ to \ gravity \\
v \ is \ the \ velocity \ of \ the \ fluid \\
\Delta h \ is \ the \ change \ into \ the \ height ."


(1)

"P_1 +\frac{1}{2}\rho v^2 +\rho gh_1 = P_2+\frac{1}{2}\rho v^2 +\rho gh_2" (velocity is same)


"P_2= P_1 + \rho gh_1 - \rho gh_2"

"= P_1 + \rho g(h_1-h_2)"

"=P_1 + \rho g(\Delta h)"

"= 3 \times 10^5 + 1000 \times 9.8 \times 2.5"

"\boxed{P = 3.24 \times 10 ^5 N/m^2}"



(2)


Fron the rate law

"Q = Av"


by given information

we know that

"Q = A_1v_1 \\
v_1 = \frac Q A_1 \\
\\"


= "\frac Q {\pi r_1^2}"

"= \frac{0.750\times 10^{-3}}{\pi \times 0.015^2}"

"= 1.06"

"\boxed{V_1 = 1.06 m/s}"


now

"Q = A_2v_2 \\
v_2 = \frac Q A_2 \\
\\"


= "\frac Q {\pi r_2^2}"

"= \frac{0.750\times 10^{-3}}{\pi \times 0.02^2}"

"= 0.597"

"\boxed{V_2 =0.597m/s}"



now

using

"P_2 = P_1+\frac{1}{2}\rho (v_1-v_2) + \rho g (h_1-h_2)\\"

"= 3\times 10^5 - \frac{1}{2}\times 1000(1.06^2-0.597^2) +1000\times 9.8 \times 2.5"

"= 2.75\times 10^5"

"\boxed{P_2 = 2.75 \times 10^5 N/m^2}"